OpenStudy (anonymous):

Fill in the missing values to make the equations true. log_4(7)+log_4(5)=log_4(?) log_8(11)-log_8(?)=log_8(11/5) 2log_9(3)=log_9(?) is the first one 12? is the second one 5? the third 3/2?

OpenStudy (anonymous):

I think adding logs means multiplying their arguments inside the log. In that case it wouldn't be 12 for the first equation but...?

OpenStudy (anonymous):

The second one is right: subtracting logs is like dividing the arguments.

OpenStudy (anonymous):

the first one you multiply:) thanks

OpenStudy (anonymous):

And in the third case: multiplication outside the log becomes potentiacion inside it. So rather than saying 3 divided by 2 you'd say three ... 2. Let me know if that doesn't help...

OpenStudy (anonymous):

Oups, sorry for my horrible orthography. :-D

OpenStudy (campbell_st):

(1) log(a) + log(b) = log(a x b) the missing part is 7 x5 or 35 (2) log(a - log(b) = log(a/b) the missing part is 5 (3) log(a^b) = blog(a) the missing part 3^2 or 9 which would give an answer of 1 (4)

OpenStudy (anonymous):

I'm remembering this like so: If you make a list of operations with increasing "strength" or "power" then addition/subtraction would be the "weakest", multiplication would be "middle" and exponentiation would be "strongest". When you go from outside a log to inside a log then you have to increase the "strength" one step: addition/subtraction outside log becomes multiplication/division inside. And multiplication/division outside becomes exponentiation inside. If this seems just complicated then forget it. Just my personal way to remember it.